Calculate the mean score of 40 students in a class text

Scores. F x

0_10. 4 5

10_20. 6 15

20_30. 8 25

30_40. 10 35

40_50. 12 45


1
Expert's answer
2021-11-22T09:22:19-0500

scores F X FX

0 - 10 4 5 20

10 - 20 6 15 90

20 - 30 8 25 200

30 - 40 10 35 350

40 - 50 12 45 540

Total 40 1200


Fx is given by F* X


Mean = "\sum" fx / "\sum"f

=1200/40

= 30